Circular No. 5977

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

GRO J2014+34 M. H. Finger, Universities Space Research Association; M. Stollberg

and G. N. Pendleton, University of Alabama, Huntsville; R. B. Wilson, Marshall Space Flight Center, NASA; and D. Chakrabarty, J. Chiu, and T. A. Prince, California Institute of Technology, report for the Compton Gamma-Ray Observatory BATSE team: “Pulsed hard x-ray emission is being seen from a previously uncatalogued object in the Cygnus-Vulpecula region. The current best-fit source position is α = 20^{h}14^{m}, δ = +34^{o}.5 (equinox 2000.0), with an estimated 68-percent-confidence region of radius 8^{o}. Work is continuing to verify this position estimate. The emission from the pulsar was first detectable on Apr. 6 and reached peak intensity on Apr.

16. The source is detected at energies from 20 up to at least 70 keV. The barycentric pulse period on Apr. 14.0 UT was 18.7021 ± 0.0001 s, with the period increasing at a rate of (6.0 ± 0.7) × 10^{−9}. The pulse profile averaged over Apr. 14–16 has a main peak with a FWHM of ∼ 0.5 in phase, with a 0.15-HWHM ‘shoulder’ present on the falling edge. The 20- to 40-keV pulsed flux on Apr. 16 was ∼ 50 mCrab, with the spectrum well fitted by an optically thin thermal bremsstrahlung model F (E) = (A/E)exp(−E/kT), with temperature kT = 23 ± 6 keV and F (40 keV) = 1.88 ± 0.24 × 10^{−4} photon cm^{−2} s^{−1} keV^{−1}.”

GRO J1008–57 M. J. Coe, C. Everall, and P. D. Roche, Southampton University; J.

Fabregat, Valencia University; D. A. H. Buckley, South African Astronomical Observatory; and R. C. Smith, Cerro Tololo Interamerican Observatory, report: “We have found a probable optical counterpart to the transient pulsar GRO J1008–57 (cf. IAUC 5836, 5838, 5851, 5877). A U.K. Schmidt red plate taken on 1993 July 28 (during discovery outburst) reveals two objects within the ROSAT error circle, the brightest at α = 10^{h}09^{m}46^{s}.9, δ = −58^{o}17′35′′ (equinox 2000.0). Spectroscopic and photometric observations from the SAAO 1.9-m telescope (Dec. 1993), CTIO Schmidt telescope (Jan. 1994), and the 3.9-m Anglo-Australian Telescope (Mar. 1994) reveal this brightest star as the only Hα emitter in the region, with strong infrared emission (J = 11.2, H = 10.5, K = 9.9). This suggests that the system is another Be or supergiant x-ray binary, as indicated by the x-ray characteristics.”

1994 April 20 Daniel W. E. Green